# Landsat30-AU: A Vision-Language Dataset for Australian Landsat Imagery

### v2

@@ Abstract
Vision language models (VLMs) that enable natural language interaction with satellite imagery can democratize Earth observation by accelerating expert workflows, making data accessible to non-specialists, and enabling planet-scale automation. However, existing datasets focus mainly on short-term, high-resolution imagery from a limited number of satellites, overlooking low-resolution, multi-satellite, long-term archives, such as Landsat, that are essential for affordable and bias-robust global monitoring. We address this gap with Landsat30-AU, a large-scale vision-language dataset built from 30-meter resolution imagery collected by four Landsat satellites (5, 7, 8, and 9) over Australia, spanning more than 36 years. The dataset includes two components: Landsat30-AU-Cap, containing $196,262$ image-caption pairs, and Landsat30-AU-VQA, comprising 17,725 human-verified visual question answering (VQA) samples across eight remote sensing domains. Both datasets are curated through a bootstrapped pipeline that leverages generic VLMs with iterative refinement and human verification to ensure quality. Our evaluation of eight VLMs on our benchmark reveals that off-the-shelf models struggle to understand satellite imagery. The open-source remote-sensing VLM EarthDial achieves only 0.07 SPIDEr in captioning and a VQA accuracy of 0.48, highlighting the limitations of current approaches. Encouragingly, lightweight fine-tuning of Qwen2.5-VL-7B on Landsat30-AU improves captioning performance from 0.11 to 0.31 SPIDEr and boosts VQA accuracy from 0.74 to 0.87. Code and data are available at https://github.com/papersubmit1/landsat30-au.
@@ 1 Introduction
For over fifty years, the Landsat program has provided a globally consistent, open-access archive of optical satellite imagery at 30-meter ground-sample distance (GSD) (Wulder et al. 2022). Since 1972, eight Landsat satellites have been launched, each equipped with distinct sensors and band configurations, resulting in varying appearances of standard red-green-blue composites across missions (U.S. Geological Survey 2025). The upcoming Landsat Next series will significantly increase daily acquisition volume, from 900 GB (750 scenes) to 8.2 TB (2,220 scenes), through expanded spectral coverage and improved sensor technology (U.S. Geological Survey 2024; NASA Landsat Science 2024). Meanwhile, vision-language models (VLMs) have shown impressive capabilities in managing and interpreting large-scale Earth observation data, especially with high-resolution sources such as Sentinel-2 imagery (Kuckreja et al. 2024; Zhang et al. 2024; Bazi et al. 2024; Yuan et al. 2024). These trends present a timely opportunity: VLMs could serve as natural-language interfaces for long-term, cost-effective, and planet-scale analysis using the growing Landsat archive.
@@
Progress is nevertheless constrained by the absence of large-scale image-text corpora that match Landsat’s unique regime. Most existing remote-sensing datasets (i) focus on sub-meter commercial imagery, which encourages captions centered on fine-grained objects, such as cars, rooftops, or road markings, that are invisible at 30 m resolution, and often come with restrictive licensing costs that hinder global-scale applications (Qu et al. 2016; Ge et al. 2025); (ii) cover only one or two Landsat satellites, preventing VLMs from learning the radiometric and band-layout differences that span the full eight-mission Landsat program, and thus limiting their robustness to sensor shifts; and (iii) include Landsat imagery with only a short temporal span, depriving models of exposure to long-term seasonal patterns, land-cover change, and climate-driven dynamics critical for temporal generalization. For example, the datasets that do incorporate Landsat imagery fall short: EarthDial includes 1.6 million image patches, but only from Landsat 8 (Soni et al. 2025), while SSL4EO-L provides five million multi-temporal patches across several missions, yet lacks the associated textual supervision necessary for vision-language alignment (Stewart et al. 2023). As a result, there is still no dataset that delivers the long-term, multi-satellite, and resolution-aware supervision needed to develop VLMs for scalable and bias-robust Earth monitoring.
@@
Generating high-quality text annotations for remote sensing images also remains a significant challenge. Manual captioning by domain experts ensures high accuracy (Qu et al. 2016; Zhan, Xiong, and Yuan 2023) but does not scale. Crowdsourcing or automatic alternatives, such as OpenStreetMap (OSM) (OpenStreetMap contributors 2025) tags or web alt-text (Muhtar et al. 2024a; Zavras et al. 2025), offer scalability but suffer from two key issues: (i) spatial mismatch, where many labeled objects (e.g., clinic, cemetery) are too small to be resolved in 30 m Landsat imagery, and (ii) temporal misalignment, where the metadata may describe a scene years before or after the satellite image was acquired, leading to outdated associations.
@@
To address these limitations, we present Landsat30-AU, the first large-scale vision-language dataset constructed entirely from 30-meter resolution imagery captured by four Landsat missions (5, 7, 8, and 9) across Australia, spanning from 1988 to 2024. It consists of two parts: (i) Landsat30-AU-Cap, containing $196,262$ image-caption pairs, and (ii) Landsat30-AU-VQA, comprising $17,725$ human-verified visual question answering (VQA) examples covering eight common remote sensing tasks. To address the challenges of scale and label quality, we develop a semi-automatic bootstrapped pipeline that extends the methodologies of HRS-Align (Muhtar et al. 2024b) and VRSBench (Li, Ding, and Elhoseiny 2024). In this pipeline, generic VLMs generate initial drafts guided by coarse, noisy metadata-spatially and temporally aligned information such as land-cover maps and crowdsourced OSM tags. Successive VLM-assisted refinement steps polish these drafts, and human reviewers remove any text that is visually ungrounded or temporally mismatched. By pairing multi-satellite, multi-decadal Landsat scenes with reliable language supervision, Landsat30-AU provides the first solid foundation for training and evaluating VLMs aimed at affordable, long-term Earth monitoring.
@@ 2 Related Work
Large-scale image-text datasets play an important role in the development of VLMs. Pioneering VLM datasets like Flickr30k (Plummer et al. 2016) and MS COCO (Lin et al. 2015) relied on costly human annotation. The SBU Captioned Photo Dataset (Ordonez, Kulkarni, and Berg 2011) and Conceptual Captions 3M (Sharma et al. 2018) expanded the scale of VLM datasets to several million image-text pairs by using web images and their associated alt-text. Using a similar approach and adding quality control from machine learning models, VLM datasets like Conceptual 12M (Changpinyo et al. 2021), LAION-5B (Schuhmann et al. 2022), and COYO-700M (Byeon et al. 2022) further increased the scale to hundreds of millions or even billions of pairs. The success of models like CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021) demonstrated that even large-scale datasets with noisy information can significantly contribute to VLM development. Many researchers are working to improve dataset quality by using advanced VLMs, like BLIP (Li et al. 2022) and InstructBLIP (Dai et al. 2023), to refine noisy data and generate higher-quality annotations. Furthermore, models such as LLaVA (Liu et al. 2023) and MiniGPT-4 (Zhu et al. 2023) generate synthetic captions to build large-scale training datasets and reduce dataset costs.
@@
The evolution of remote sensing VLMs has mirrored that of the general VLM community. Datasets like UCM-Captions and Sydney-Captions (Lu et al. 2018) consisted of only a few hundred images with domain expert labels. To increase dataset scale, NWPU-Captions (Cheng et al. 2022) and RSICD (Lu et al. 2018) retrieved imagery and metadata from commercial map services and then used crowdsourcing to edit the captions. With imagery from open-source satellite platforms, SkyScript (Wang et al. 2023) and OpenSentinelMap (Johnson, Treible, and Crispell 2022) utilized open-source tags from free map services to create captions. However, this approach introduces temporal misalignments between static map tags and dynamic landcover. Following the success of synthetic datasets in general VLMs, remote sensing projects such as RS5M (Zhang et al. 2024), SkySenseGPT (Luo et al. 2024), ChatEarthNet (Yuan et al. 2024), Git-10M (Liu et al. 2025), and RS-LLaVA (Bazi et al. 2024) have scaled to tens of millions of image-text pairs by using prompted LLMs to synthesize instructions. Meanwhile, task-specific VQA benchmarks such as RSIVQA (Lobry et al. 2020) continue to reveal VLM weaknesses in counting, spatial reasoning, and domain transfer. Despite recent progress, the historical Landsat archive remains underutilized in VLM research. For example, the recent EarthDial (Soni et al. 2025), despite its multi-sensory approach, includes only imagery from Landsat 8. Landsat30-AU addresses this gap by providing images from four Landsat sensors that span from 1988 to 2024, thereby enabling long-term, continental-scale studies with an open-source VLM dataset.
Figure 1: Overview of the Landsat30-AU dataset construction pipeline.
@@ 3 Dataset Construction
We source atmospherically and geometrically corrected imagery from the Digital Earth Australia (DEA) Analysis Ready Data (ARD) archive (Geoscience Australia 2024). We use Bands 4 (Red), 3 (Green), and 2 (Blue) to generate over 400,000 $256\times 256$ pixel RGB tiles at 30-meter GSD.
@@
The DEA Land Cover product (Geoscience Australia 2025) provides annually updated, pixel-level classifications (e.g., artificial surfaces, natural bare, water) derived from satellite observations. We extract the dominant land cover class for six fixed spatial regions within each image: top-left, top-right, bottom-left, bottom-right, center, and entire tile. These structured region-level labels support downstream tasks such as region classification and guided captioning.
@@
Following ChatEarthNet (Yuan et al. 2024), each $256{\times}256$ tile is partitioned into six zones: top-left, top-right, bottom-left, bottom-right, center, and entire tile. For each zone, we assign the dominant land-cover class based on the DEA annual land-cover raster (Geoscience Australia 2025), using a taxonomy of coarse land-cover types (e.g., cropland, forest, water, urban).
We manually validate 2,722 such tile-region label sets and fine-tune GPT-4o (OpenAI 2024) on this task. For correctness, we use Subset Accuracy (Godbole and Sarawagi 2004). For set similarity and per-label quality, we employ the Jaccard Index, Precision, Recall, and F1-score. Ranking performance is measured with Label-Ranking Average Precision (LRAP) (Elisseeff and Weston 2001) and nDCG (Järvelin and Kekäläinen 2002). To incorporate error rates, we report (1 - Hamming Loss) and (1 - Ranking Loss), ensuring higher values are consistently better across all metrics. The fine-tuned model achieves Subset Accuracy 0.28 and Jaccard 0.63, outperforming a Qwen2.5-VL-7B (Qwen) (Bai et al. 2025) baseline (Table 1(a)).
@@
We fine-tuned GPT-4.1 (OpenAI 2025) on this seed dataset, resulting in captions with broader semantic coverage and improved factual grounding. Quantitatively, SPIDEr increased from 0.44 to 0.52, indicating better alignment with reference semantics, while 1-CHAIR-s rose from 0.44 to 0.47, reflecting fewer hallucinated object mentions. The average caption length also increased from 149 to 161 tokens, suggesting greater descriptive depth (Table 1(b)).
@@
OCA 2,129 Cloud/haze usability assessment Fig. 3(b): “Scene usable despite cloud?” Options: Fully usable, Not usable
@@
Figure 3: Landsat30-AU-VQA categories.
@@
For each image tile, the captioning model (fine-tuned GPT-4.1) first generates an Initial caption conditioned on region labels, OSM tags, and the image. We then prompt Qwen2.5-VL-7B to augment the caption with missing objects and spatial relations, resulting in an Extra version. Next, the caption reviewer module prunes hallucinated or temporally inconsistent content, producing the final Reviewed caption. To evaluate the impact of each stage, we score all three versions (Initial, Extra, and Reviewed) on a held-out reference set using BLEU-4 (Papineni et al. 2002), SPIDEr (Liu et al. 2017), and BERTScore-F1 (Zhang et al. 2020) for semantic quality, and CHAIR-s/i (Rohrbach et al. 2019) for hallucination. As shown in Table 1(b), the Reviewed captions provide the best overall balance (SPIDEr 0.517; 1-CHAIR-i 0.853), and are used throughout the dataset. This process yields 196,262 high-quality captions, which make up the Landsat30-AU-Cap dataset.
@@ 4 Landsat30-AU Dataset
This section provides an overview of the two sub-datasets that make up Landsat30-AU, along with their key statistics and a comparison with existing remote sensing vision-language corpora.
@@
We compare Landsat30-AU to six key remote sensing vision-language datasets: RSICD (Lu et al. 2018), SkyScript (Wang et al. 2023), ChatEarthNet (Yuan et al. 2024), Git-10M (Liu et al. 2025), GAIA (Zavras et al. 2025), and EarthDial (Soni et al. 2025).

### v3

@@ Abstract
Vision language models (VLMs) that enable natural language interaction with satellite imagery can democratize Earth observation by accelerating expert workflows, making data accessible to non-specialists, and enabling planet-scale automation. However, existing datasets focus mainly on short-term, high-resolution imagery from a limited number of satellites, overlooking low-resolution, multi-satellite, long-term archives, such as Landsat, that are essential for affordable and bias-robust global monitoring. We address this gap with Landsat30-AU, a large-scale vision-language dataset built from 30-meter resolution imagery collected by four Landsat satellites (5, 7, 8, and 9) over Australia, spanning more than 36 years. The dataset includes two components: Landsat30-AU-Cap, containing $196,262$ image-caption pairs, and Landsat30-AU-VQA, comprising 17,725 human-verified visual question answering (VQA) samples across eight remote sensing domains. Both datasets are curated through a bootstrapped pipeline that leverages generic VLMs with iterative refinement and human verification to ensure quality. Our evaluation of eight VLMs on our benchmark reveals that off-the-shelf models struggle to understand satellite imagery. The open-source remote-sensing VLM EarthDial achieves only 0.07 SPIDEr in captioning and a VQA accuracy of 0.48, highlighting the limitations of current approaches. Encouragingly, lightweight fine-tuning of Qwen2.5-VL-7B on Landsat30-AU improves captioning performance from 0.11 to 0.31 SPIDEr and boosts VQA accuracy from 0.74 to 0.87.
Codehttps://github.com/papersubmit1/landsat30-au
@@ 1 Introduction
For over fifty years, the Landsat program has provided a globally consistent, open-access archive of optical satellite imagery at 30-meter ground-sample distance (GSD) (WULDER2022113195). Since 1972, eight Landsat satellites have been launched, each equipped with distinct sensors and band configurations, resulting in varying appearances of standard red-green-blue composites across missions (usgs_landsat_band_design). The upcoming Landsat Next series will significantly increase daily acquisition volume, from 900 GB (750 scenes) to 8.2 TB (2,220 scenes), through expanded spectral coverage and improved sensor technology (usgs_landsat_next_news; usgs_landsat_next_data_volume_change). Meanwhile, vision-language models (VLMs) have shown impressive capabilities in managing and interpreting large-scale Earth observation data, especially with high-resolution sources such as Sentinel-2 imagery (Kuckreja_2024_CVPR; 10679571; rs16091477; yuan2024chatearthnetglobalscaleimagetextdataset). Following these trends, VLMs offer a timely opportunity as natural-language interfaces for long-term, cost-effective, planet-scale analysis using the growing Landsat archive.
@@
Progress is nevertheless constrained by the absence of large-scale image-text corpora that match Landsat’s unique regime. Most existing remote-sensing datasets (i) focus on sub-meter commercial imagery, which encourages captions centered on fine-grained objects, such as cars, rooftops, or road markings, that are invisible at 30 m resolution, and often come with restrictive licensing costs that hinder global-scale applications (7546397; ge2025rstellerscalingvisuallanguage); (ii) cover only one or two Landsat satellites, preventing VLMs from learning the radiometric and band-layout differences that span the full eight-mission Landsat program, and thus limiting their robustness to sensor shifts; and (iii) include Landsat imagery with only a short temporal span, depriving models of exposure to long-term seasonal patterns, land-cover change, and climate-driven dynamics critical for temporal generalization. For example, the datasets that do incorporate Landsat imagery fall short: EarthDial includes 1.6 million image patches, but only from Landsat 8 (soni2025earthdialturningmultisensoryearth), while SSL4EO-L provides five million multi-temporal patches across several missions, yet lacks the associated textual supervision necessary for vision-language alignment (stewart2023ssl4eoldatasetsfoundationmodels). As a result, there is still no dataset that delivers the long-term, multi-satellite, and resolution-aware supervision needed to develop VLMs for scalable and bias-robust Earth monitoring.
@@
Generating high-quality text annotations for remote sensing images also remains a significant challenge. Manual captioning by domain experts ensures high accuracy (7546397; Zhan_2023) but does not scale. Crowdsourcing or automatic alternatives, such as OpenStreetMap (OSM) (openstreetmap_community) tags or web alt-text (10.1007/978-3-031-72904-1_26; zavras2025gaiaglobalmultimodalmultiscale), offer scalability but suffer from two key issues: (i) spatial mismatch, where many labeled objects (e.g., clinic, cemetery) are too small to be resolved in 30 m Landsat imagery, and (ii) temporal misalignment, where the metadata may describe a scene years before or after the satellite image was acquired, leading to outdated associations.
@@
To address these limitations, we present Landsat30-AU, the first large-scale vision-language dataset constructed entirely from 30-meter resolution imagery captured by four Landsat missions (5, 7, 8, and 9) across Australia, spanning from 1988 to 2024. It consists of two parts: (i) Landsat30-AU-Cap, containing $196,262$ image-caption pairs, and (ii) Landsat30-AU-VQA, comprising $17,725$ human-verified visual question answering (VQA) examples covering eight common remote sensing tasks. To address the challenges of scale and label quality, we develop a semi-automatic bootstrapped pipeline that extends the methodologies of LHRS-Align (10.1007/978-3-031-72904-1_26) and VRSBench (li2024vrsbenchversatilevisionlanguagebenchmark). In this pipeline, generic VLMs generate initial drafts guided by coarse, noisy metadata-spatially and temporally aligned information such as land-cover maps and crowdsourced OSM tags. Successive VLM-assisted refinement steps polish these drafts, and human reviewers remove any text that is visually ungrounded or temporally mismatched. By pairing multi-satellite, multi-decadal Landsat scenes with reliable language supervision, Landsat30-AU provides the first solid foundation for training and evaluating VLMs aimed at affordable, long-term Earth monitoring.
@@ 2 Related Work
Large-scale image-text datasets play an important role in the development of VLMs. Pioneering VLM datasets like Flickr30k (DBLP:journals/corr/PlummerWCCHL15) and MS COCO (DBLP:journals/corr/LinMBHPRDZ14) relied on costly human annotation. The SBU Captioned Photo Dataset (NIPS2011_5dd9db5e) and Conceptual Captions 3M (sharma2018conceptual) expanded the scale of VLM datasets to several million image-text pairs by using web images and their associated alt-text. Using a similar approach and adding quality control from machine learning models, VLM datasets like Conceptual 12M (DBLP:journals/corr/abs-2102-08981), LAION-5B (schuhmann2022laion5bopenlargescaledataset), and COYO-700M (kakaobrain2022coyo-700m) further increased the scale to hundreds of millions or even billions of pairs. The success of models like CLIP (DBLP:journals/corr/abs-2103-00020) and ALIGN (DBLP:journals/corr/abs-2102-05918) demonstrated that even large-scale datasets with noisy information can significantly contribute to VLM development. Many researchers are working to improve dataset quality by using advanced VLMs, like BLIP (DBLP:journals/corr/abs-2201-12086) and InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels), to refine noisy data and generate higher-quality annotations. Furthermore, models such as LLaVA (liu2023visualinstructiontuning) and MiniGPT-4 (zhu2023minigpt4enhancingvisionlanguageunderstanding) generate synthetic captions to build large-scale training datasets and reduce dataset costs.
@@
The evolution of remote sensing VLMs has mirrored that of the general VLM community. Datasets like UCM-Captions and Sydney-Captions (DBLP:journals/corr/abs-1712-07835) consisted of only a few hundred images with domain expert labels. To increase dataset scale, NWPU-Captions (9866055) and RSICD (rosario2023satellitecaptioninglargelanguage) retrieved imagery and metadata from commercial map services and then used crowdsourcing to edit the captions. With imagery from open-source satellite platforms, SkyScript (wang2023skyscriptlargesemanticallydiverse) and OpenSentinelMap (9856983) utilized open-source tags from free map services to create captions. However, this approach introduces temporal misalignments between static map tags and dynamic landcover. Following the success of synthetic datasets in general VLMs, remote sensing projects such as RS5M (10679571), SkySenseGPT (luo2024skysensegptfinegrainedinstructiontuning), ChatEarthNet (yuan2024chatearthnetglobalscaleimagetextdataset), Git-10M (10988859), and RS-LLaVA (rs16091477) have scaled to tens of millions of image-text pairs by using prompted LLMs to synthesize instructions. Meanwhile, task-specific VQA benchmarks such as RSIVQA (DBLP:journals/corr/abs-2003-07333) continue to reveal VLM weaknesses in counting, spatial reasoning, and domain transfer. Despite recent progress, the historical Landsat archive remains underutilized in VLM research. For example, the recent EarthDial (soni2025earthdialturningmultisensoryearth), despite its multi-sensory approach, includes only imagery from Landsat 8. Landsat30-AU addresses this gap by providing images from four Landsat sensors that span from 1988 to 2024, thereby enabling long-term, continental-scale studies with an open-source VLM dataset.
Figure 1: Overview of the Landsat30-AU dataset construction pipeline. Stage 1: Sources Landsat imagery and collects metadata. Stage 2: Adapts VLMs into specialized modules for region classification, caption generation, and review. Stage 3: Produces large-scale annotations via iterative VLM refinement and human verification.
@@ 3 Dataset Construction
We source atmospherically and geometrically corrected imagery from the Digital Earth Australia (DEA) Analysis Ready Data (ARD) archive (ga_analysis_ready_data). We use Bands 4 (Red), 3 (Green), and 2 (Blue) to generate over 400,000 $256\times 256$ pixel RGB tiles at 30-meter GSD.
@@
The DEA Land Cover product (ga_dea_landcover_v2) provides annually updated, pixel-level classifications (e.g., artificial surfaces, natural bare, water) derived from satellite observations. We extract the dominant land cover class for six fixed spatial regions within each image: top-left, top-right, bottom-left, bottom-right, center, and entire tile. These structured region-level labels support downstream tasks such as region classification and guided captioning.
@@
Following ChatEarthNet (yuan2024chatearthnetglobalscaleimagetextdataset), each $256{\times}256$ tile is partitioned into six zones: top-left, top-right, bottom-left, bottom-right, center, and entire tile. For each zone, we assign the dominant land-cover class based on the DEA annual land-cover raster (ga_dea_landcover_v2), using a taxonomy of coarse land-cover types (e.g., cropland, forest, water, urban).
We manually validate 2,722 such tile-region label sets and fine-tune GPT-4o (gpt-4o-2024-08-06) from OpenAI (openai2024gpt4ocard) on this task. For correctness, we use Subset Accuracy (10.1007/978-3-540-24775-3_5). For set similarity and per-label quality, we employ the Jaccard Index, Precision, Recall, and F1-score. Ranking performance is measured with Label-Ranking Average Precision (LRAP) (10.5555/2980539.2980628) and nDCG (jarvelin2002cumulated). To incorporate error rates, we report (1 - Hamming Loss) and (1 - Ranking Loss), ensuring higher values are consistently better across all metrics. The fine-tuned model achieves Subset Accuracy 0.28 and Jaccard 0.63, outperforming a Qwen2.5-VL-7B (Qwen) (bai2025qwen25vltechnicalreport) baseline (Table 1(a)).
@@
We fine-tuned GPT-4.1 (gpt-4.1-2025-04-14) from OpenAI (openai_gpt41) on this seed dataset, resulting in captions with broader semantic coverage and improved factual grounding. Quantitatively, SPIDEr increased from 0.44 to 0.52, indicating better alignment with reference semantics, while 1-CHAIR-s rose from 0.44 to 0.47, reflecting fewer hallucinated object mentions. The average caption length also increased from 149 to 161 tokens, suggesting greater descriptive depth (Table 1(b)).
@@
COA 2,129 Cloud/haze usability assessment Fig. 3(b): “Scene usable despite cloud?” Options: Fully usable, Not usable
@@
Figure 3: Landsat30-AU-VQA categories. Representative 30-meter Landsat imagery illustrating the visual characteristics of each of the eight VQA domains.
@@
For each image tile, the captioning model (fine-tuned GPT-4.1) first generates an Initial caption conditioned on region labels, OSM tags, and the image. We then prompt Qwen2.5-VL-7B to augment the caption with missing objects and spatial relations, resulting in an Extra version. Next, the caption reviewer module prunes hallucinated or temporally inconsistent content, producing the final Reviewed caption. To evaluate the impact of each stage, we score all three versions (Initial, Extra, and Reviewed) on a held-out reference set using BLEU-4 (10.3115/1073083.1073135), SPIDEr (DBLP:journals/corr/LiuZYG016), and BERTScore-F1 (DBLP:journals/corr/abs-1904-09675) for semantic quality, and CHAIR-s/i (DBLP:journals/corr/abs-1809-02156) for hallucination. As shown in Table 1(b), the Reviewed captions provide the best overall balance (SPIDEr 0.517; 1-CHAIR-i 0.853), and are used throughout the dataset. This process yields 196,262 high-quality captions, which make up the Landsat30-AU-Cap dataset.
@@ 4 Landsat30-AU Dataset
This section details the two Landsat30-AU sub-datasets, providing key statistics and a comparison with existing remote sensing vision-language corpora.
@@
We compare Landsat30-AU to six key remote sensing vision-language datasets: RSICD (rosario2023satellitecaptioninglargelanguage), SkyScript (wang2023skyscriptlargesemanticallydiverse), ChatEarthNet (yuan2024chatearthnetglobalscaleimagetextdataset), Git-10M (10988859), GAIA (zavras2025gaiaglobalmultimodalmultiscale), and EarthDial (soni2025earthdialturningmultisensoryearth).